OpenStudy (anonymous):

Please help... Determine the measure of side c, given: a=3m b=4m B=41deg

OpenStudy (aravindg):

tried cosine rule ?

OpenStudy (anonymous):

im not sure how to do this at all....

OpenStudy (aravindg):

do u knw cosine rule ?

OpenStudy (anonymous):

i have just started learning about it

OpenStudy (aravindg):

ok \[\cos B=\dfrac{a^2+c^2-b^2}{2ac} \]

OpenStudy (aravindg):

now solve for b

OpenStudy (anonymous):

yea good cosine rule,,my answer is b=2.44 (approx.)

OpenStudy (dan815):

u could use sin law too i think

OpenStudy (anonymous):

sorry 1.22

OpenStudy (aravindg):

@dan815 sin law wont work cause we need two angles :/

OpenStudy (dan815):

sin <b/b=sin<a=a sin<(180-(a+b))/c and solve for C

OpenStudy (dan815):

use my equation < means angle

OpenStudy (anonymous):

actually my options are: 1.6m 4.5m 0.1m 5.7m I thought it was 1.6m?

OpenStudy (anonymous):

am I wrong?

OpenStudy (anonymous):

actually it was 5.7 after I figured it out.. Thanks though

OpenStudy (dan815):

did u use my formula!

OpenStudy (anonymous):

yea approx.

OpenStudy (anonymous):

yes thank you

OpenStudy (dan815):

np :)

OpenStudy (dan815):

if u remember that you need atleast 3 pieces of information to solve a triangle and that if the angle given is the angle of the opposite side then you can always use sin law if they give u like side a and b but give you angle C then you have to use cos law if its not a right angle triangle